OpenStudy (anonymous):

The total rainfall in Forks, Washington, this October was 8 3/8 in. In September, 60% less rain fell than in October. How much rain fell in September in Forks, Washington? Express your answer as a mixed number in simplest form.

OpenStudy (anonymous):

medal

OpenStudy (anonymous):

@mariod5

OpenStudy (anonymous):

\[\frac{ 201 }{ 40 }\]

OpenStudy (anonymous):

in smalest form

OpenStudy (anonymous):

that is the smallest form

OpenStudy (zpupster):

October was 8 3/8 in September, 60% less rain 100-60= 40 percent of(mult) October rain so .40 * 8 3/8=

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

do i get a medal?

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

3.35 is the answer to the other guys question

OpenStudy (anonymous):

has to be a fraction thou

OpenStudy (anonymous):

7/20

OpenStudy (anonymous):

r u sure

OpenStudy (anonymous):

3 7/20

OpenStudy (anonymous):

r u sure

OpenStudy (anonymous):

thats the answer to @zpupster question

OpenStudy (anonymous):

and yes im sure

OpenStudy (anonymous):

ok but is it right